# Honeybee: Locality-enhanced Projector for Multimodal LLM

### v1

@@ 1 Introduction
Notably, recent MLLMs prefer abstractors to linear projectors; this is primarily due to their flexibility in handling the number of resulting visual tokens, thus offering versatile design options for achieving a preferable balance between efficiency and effectiveness. However, according to our observation in Fig. 3, it turns out that the abstractors face challenges when learning tasks oriented towards spatial understanding compared to the linear projectors. This difficulty arises from the absence of a locality-aware design during the abstraction process, leading it to primarily abstract visual information from a few regions rather than retaining information about all regions, thus losing finer details crucial for spatial comprehension. In contrast, linear projectors excel at preserving the local context of visual features via one-to-one transformation. This strong preservation of locality allows effective spatial understanding.
Motivated by this, we propose novel locality-enhanced abstractors as the projector, which exhibit a more favorable balance between performance (by locality preservation) and efficiency (by abstraction capability) as presented in Fig. 1. To be specific, we introduce two locality-enhanced abstractors, C-Abstractor and D-Abstractor, by employing two powerful operations in locality modeling—convolution and deformable attention. Such injection of locality-aware design into abstractors not only promotes the overall performance improvement of MLLMs in handling intricate visual information but also capitalizes on computational efficiency during the subsequent response generation phase of LLMs.
@@
We identify two important projector properties: 1) locality preservation of visual features and 2) flexibility to manage the number of visual tokens, and propose locality-enhanced abstractors that achieve the best of both worlds.
@@ 3.1 Overview
Generally, the goal of Multimodal Large Language Models (MLLMs) is to learn a model that can produce instruction-following responses for the given multimodal inputs. In this paper, we consider images as additional modality inputs to MLLMs. Thus, the language model becomes a receiver of both visual and text (instruction) tokens while generating text responses in an autoregressive manner. Formally, a multimodal input consists of two types of tokens: image tokens and text tokens . Then, the language model predicts the response conditioned on the multimodal input where means the number of tokens in the response. Therefore, the response is predicted by
@@ 3.2.2 Architecture
In deep learning, convolution has been the most successful architecture for modeling local context [24, 50, 48]. Thus, we design Convolutional Abstractor, C-Abstractor, for effective local context modeling. Fig. 4a depicts the entire architecture, comprising ResNet blocks [50] followed by adaptive average pooling and another ResNet blocks. This design allows to abstract visual features to any squared number of visual tokens, and even project to more visual tokens than the original number of visual features. We also tested several variants [36, 48], but ResNet [50] shows the best performance. Further details are provided in Appendix B.
@@ 3.3 Training
We train Honeybee in the two-stage pipeline. In the first stage, we freeze the vision encoder and LLM, focusing on training the proposed locality-enhanced abstractor. In the second stage, we train both the projector and LLM to enhance deeper visual understanding and generation abilities.
@@
After the pre-training of the projector for vision-language alignment, in the second stage, we jointly train the projector and LLM to enhance instruction-following capabilities and achieve a more profound visual understanding. For instruction-following, we utilize two GPT-assisted instruction-following datasets, LLaVA [33] and ShareGPT [10]. In addition, to enhance visual understanding, we instructize a wide range of existing datasets, as listed in Table 2. Specifically, our approach includes: 1) employing a range of tasks such as VQA [16, 20, 41, 30], multiple-choice VQA [38, 47], captioning [5, 26], and referring expression comprehension (visual grounding and grounded captioning) [23, 21, 40, 56]; 2) using multiple datasets for each task; 3) applying a fine-grained but single template for each dataset. Detailed examples and descriptions are in Appendix E. We thoroughly explore template-based instruction formatting strategies and the utilization of multifaceted datasets in Section 4.
@@ 4 Hidden Recipe for Visual Instruction Tuning
In Section 3, we examine the limitations of current projectors and propose methods for enhancing locality. However, a clear recipe for training cutting-edge Multimodal LLMs (MLLMs) remains unclear. While it is widely known that training MLLMs using existing datasets as instruction tuning by template-based instruction formatting is beneficial [2, 32, 11], the details of the instructization process are still underexplored—questions persist regarding dataset selection, utilization, and combination strategies. In this section, we aim to clarify these aspects via the five research questions: () To what extent does each dataset contribute to the performance of specific tasks? () What is an effective balancing strategy between diverse datasets? () What is the appropriate granularity for the templates? () How significant is the diversity of the templates? () Do conversation-like multi-turn templates provide additional benefits?
@@ 5.1 Evaluation Setting
5.2 Implementation Details
We employ 7B and 13B Vicuna-v1.5 [10] as the language model. We leverage the pre-trained CLIP ViT-L/14 [45] with resolutions of 224 and 336 for 7B- and 13B-LLM, respectively. We use the features from the second-last layer of CLIP instead of the last layer. Any image indicator tokens, e.g., special tokens enclosing visual tokens, are not used. We train the entire LLM instead of parameter-efficient fine-tuning. The long (200k pre-training, 10k instruction tuning) and short (50k pre-training, 4k instruction tuning) training schedules are used for final model comparisons and detailed analyses, respectively. The short schedule is applied with Vicuna-7B, CLIP ViT-L/14, and C-Abstractor with 224 resolution and =144 visual tokens unless stated otherwise. See Appendix C for more details.
| Projector | # Visual Tokens | MME | MMBench | SEED-Bench | Avg | |||
| POS | SR | OL | PR | SR | IL | |||
| Resampler | 144 | 75.0 | 22.2 | 43.2 | 62.5 | 47.5 | 50.6 | 43.9 |
| Linear | 256 | 140.0 | 24.4 | 40.7 | 70.8 | 48.9 | 60.9 | 52.6 |
| C-Abstractor | 144 | 135.0 | 24.4 | 54.3 | 66.7 | 49.0 | 58.8 | 53.5 |
| D-Abstractor | 144 | 138.3 | 24.4 | 45.7 | 70.8 | 49.3 | 57.8 | 52.9 |
@@ 5.2 Implementation Details
5.3 Analysis on Locality-Enhanced Projector
Spatial understanding capability.
To investigate the impact of local context preservation, we compare the spatial understanding capability on six tasks from MME, MMBench, and SEED-Bench. Table 3 summarizes the results; notably, the resampler, without consideration of local context preservation, shows poor performance. Locality-aware modeling in our projectors dramatically improves the spatial understanding capability compared to the resampler. Also, our projectors show comparable or improved performance over the linear projector with even better efficiency.
Performance-efficiency balance.
Fig. 1 presents a comparison in terms of performance vs. efficiency while varying the number of visual tokens. The linear projector cannot offer flexibility due to its one-to-one conversion. Resampler and C-Abstractor provide flexible design capabilities, allowing us to customize the model to meet different requirements with a preferable balance between efficiency and effectiveness. While the resampler suffers from limited performances, our method using 144 or 256 visual tokens performs better than the linear counterpart.
@@ 5.4 Hidden Recipe for Visual Instruction Tuning
Table 4 shows our comprehensive ablation study to identify the individual impact of datasets on various multimodal benchmarks. First, we investigate the significance of dataset diversity within each task type, by training Honeybee on the single dataset from each task type. This reveals an overall performance drop (D1 vs. D2), underscoring the importance of the diversity of datasets within each task type. Subsequent analysis investigates the impact of each task type by sequentially excluding specific tasks (D1 vs. D3-6), with the exclusion of open-ended VQA tasks notably decreasing benchmark scores. This suggests that diverse multimodal knowledge of these datasets enriches MLLM knowledge across various dimensions. Meanwhile, excluding multiple-choice VQA tasks significantly affects scores in benchmarks such as MMB and SEED, highlighting their role in aligning response patterns. The absence of captioning data particularly reduces LLaVA scores to 59.8, implying LLaVA benchmark’s preference for narrative and descriptive responses, and the importance of captioning data in training. Lastly, the exclusion of visual or text instruction-following datasets (D1 vs. D7-10) significantly impacts LLaVA, reiterating the necessity of these datasets for instruction-following ability. In summary, these experiments show the importance of task diversity in training MLLM, encompassing a variety of task types and datasets within each task type.
@@
| Method | LLM | Projector | Vision Enc. | Res. | MMB | MME | MME | SEED | LLaVA |
@@
Apart from dataset and instructization strategies, training recipes also incorporate several subtle yet crucial design choices, including the selection of features in vision encoder, LLMs, LLM training techniques, image indicators, pre-training and instruction tuning iterations. These recipes are detailed in Appendix D.
@@ 5.5 Putting It Altogether
In Table 6, we compare our Honeybee, optimized as previously discussed, with other state-of-the-art MLLMs. Honeybee outperforms comparable 7B-scale MLLMs in all benchmarks, except for SEED. It is worth noting that competing methods like Qwen-VL [2] and LLaVA-1.5 [32] use larger vision encoders (e.g., ViT-bigG for Qwen-VL) or larger images (448 and 336) with increased visual tokens (256 and 576), while Honeybee employs ViT-L/14 with 224 resolution and 144 visual tokens. Given the focus on the detailed visual understanding of SEED (See Appendix F), larger images or more visual tokens can be beneficial. With increased visual tokens (144 to 256), Honeybee achieves the best score in SEED (65.5) with 7B-scale LLM, as shown in Table 7. With 13B-scale LLMs, Honeybee surpasses all previous methods in every benchmark. The detailed scores are available in Sec. G.1.
5.6 Additional Results
@@
In our final 7B and 13B models, we use 144 and 256 visual tokens () respectively, balancing efficiency and performance. As indicated in Fig. 1 and Appendix A, increasing consistently improves performance. Our experiments, aligning in Honeybee with the linear projector (Table 7), show performance enhancement at the cost of efficiency. Additional comparisons with previous methods are in Sec. G.2.
@@
ScienceQA [38] evaluation results are presented in Sec. G.3. Remarkably, without any specialized fine-tuning, our generalist Honeybee achieves state-of-the-art performance (94.39), outperforming specialist models such as MM-CoT (91.68) [62] and LLaVA+GPT-4 (92.53) [33].
Qualitative examples are provided in Sec. H.2.
@@ 6 Conclusion
The advent of visual instruction tuning has brought remarkable advances in MLLMs. Despite these strides, areas such as projector design and the approach in handling multifaceted data with instructization processes remain underexplored or unclear. Addressing these gaps, we identify the desirable but overlooked projector property, i.e., locality preservation, and propose the locality-enhanced projector offering a preferable balance between performance and efficiency. In addition, we provide extensive experiments to identify the impact of individual design choices in handling multifaceted instruction data, unveiling hidden recipes for high-performing MLLM development. Finally, our proposed MLLM, Honeybee, remarkably outperforms previous state-of-the-art methods across various benchmarks.
@@ B.1 Linear Projector
In the recent study, LLaVA (v1.5) [32] utilizes a 2-layer MLP instead of a single linear projection for enhancing the vision-language connector’s representation power. This approach led to an investigation of how varying the number of MLP layers impacts overall performance. As shown in Table 9, the 2-layer MLP-based projector marginally improves the overall performance compared to the linear projector. However, we observe a slight performance drop when further increasing the number of MLP layers (i.e., 6-layer MLP). We note that our C- and D-Abstractors achieve better or comparable benchmark scores while using fewer visual tokens, indicating our abstractors’ superiority regarding the balance of efficiency and performance.
@@ B.2 Resampler
As described in the main text, our design focuses on two principles: 1) flexibility in visual token counts, which is the key factor to the efficiency of MLLM, and 2) preservation of local context, which is critical for spatial understanding. Our first try is augmenting visual features with positional embeddings in the resampler framework, but it does not yield notable improvements (See Resampler in Table 9). This leads us to design two novel abstractors, C-Abstractor and D-Abstractor.
@@ Appendix C Implementation Details
The detailed hyperparameters (HPs) are summarized in Table 11. Additionally, we utilize total six blocks in both C-Abstractor and D-Abstractor (i.e., for C-Abstractor and for D-Abstractor in Fig. 4). We use a single node with A100 80GB 8, employing deepspeed zero-2 [46] and flash-attention v2 [12] for all experiments, except for the long schedule pre-training where we use multinode setups.
l—l—m12cm
@@
A-OKVQA Answer with the option’s letter from the given choices directly. {question} There are several options: {option} AI: {answer}

### v2

@@ 1 Introduction
Notably, recent MLLMs prefer abstractors (e.g., resampler, Q-former) to linear projectors; this is primarily due to their flexibility in handling the number of resulting visual tokens, thus offering versatile design options for achieving a preferable balance between efficiency and performance. However, as shown in Fig. 3, the abstractors face more difficulties in learning spatial understanding tasks compared to the linear projectors. This difficulty stems from the abstraction process lacking a locality-aware design, which causes it to primarily focus on a few regions, leading to a loss of finer details essential for spatial comprehension. In contrast, linear projectors excel at preserving all local contexts of visual features via one-to-one transformation. This strong preservation of locality allows effective spatial understanding.
Motivated by this, we propose novel locality-enhanced projectors, which exhibit a more favorable balance between performance (by locality preservation) and efficiency (by abstraction capability) as presented in Fig. 1. To be specific, we introduce two locality-enhanced projectors by employing two powerful operations in locality modeling—convolution and deformable attention. Such injection of locality-aware design into the abstraction process not only promotes the overall performance improvement of MLLMs in handling intricate visual information but also capitalizes on computational efficiency during the subsequent response generation phase of LLMs.
@@
We identify two crucial properties of projector, 1) locality preservation of visual features and 2) flexibility to manage the number of visual tokens, and propose locality-enhanced abstractors to achieve the best of both worlds.
@@ 3.1 Overview
Generally, the goal of Multimodal Large Language Models (MLLMs) is to learn a model that can produce instruction-following responses for the given multimodal inputs. In this paper, we consider images as an additional modality input to MLLMs. Thus, the language model becomes a receiver of both visual and text (instruction) tokens while generating text responses in an autoregressive manner. Formally, a multimodal input consists of two types of tokens: image tokens and text tokens . Then, the language model predicts the response conditioned on the multimodal input where means the number of tokens in the response. Therefore, the response is predicted by
@@ 3.2.2 Architecture
In deep learning, convolution has been the most successful architecture for modeling local context [24, 51, 49]. Thus, we design Convolutional Abstractor, C-Abstractor, for effective local context modeling. Fig. 4a depicts the entire architecture, comprising ResNet blocks [51] followed by adaptive average pooling and another ResNet blocks. This design allows to abstract visual features to any squared number of visual tokens, and even project to more visual tokens than the original number of visual features. We also tested several variants [37, 49] in Appendix B, but ResNet [51] shows the best performance.
@@ 3.3 Training
We train Honeybee in the two-stage pipeline. In the first stage, we freeze the vision encoder and LLM, focusing on training the proposed locality-enhanced projector. In the second stage, we train both the projector and LLM to enhance deeper visual understanding and generation abilities.
@@
After the pre-training of the projector for vision-language alignment, in the second stage, we jointly train the projector and LLM to enhance instruction-following capabilities and achieve a more profound visual understanding. For instruction-following, we utilize two GPT-assisted instruction-following datasets, LLaVA [34] and ShareGPT [10]. In addition, to enhance visual understanding, we instructize a wide range of existing datasets using templates, as listed in Table 2. Specifically, our approach includes: 1) employing a range of tasks such as open-ended VQA [16, 20, 42, 31], multiple-choice VQA [39, 48], captioning [5, 26], and referring expression comprehension (visual grounding and grounded captioning) [23, 21, 41, 57]; 2) using multiple datasets for each task; 3) applying a fine-grained but single template for each dataset. Detailed examples and descriptions are in Appendix E. We thoroughly explore template-based instructization strategies and the utilization of multifaceted datasets in Section 4.
@@ 4 Hidden Recipe for Visual Instruction Tuning
In Section 3, we examine the limitations of current projectors and propose methods for enhancing locality. However, a clear recipe for training cutting-edge Multimodal LLMs (MLLMs) remains unclear. While it is widely known that instruction tuning using existing datasets with the template-based instructization is beneficial [2, 33, 11], the details of the instructization process are still underexplored—questions persist regarding dataset selection, utilization, and combination strategies. In this section, we aim to clarify these aspects via following the five research questions: () To what extent does each dataset contribute to the performance of specific tasks? () What is an effective balancing strategy between diverse datasets? () What is the appropriate granularity for the templates? () How significant is the diversity of the templates? () Do conversation-like multi-turn templates provide additional benefits?
@@ 5.1 Settings
Implementation details.
We use 7B and 13B Vicuna-v1.5 [10] as LLM. We leverage the pre-trained CLIP ViT-L/14 [46] with 224 and 336 resolutions for 7B- and 13B-LLM, respectively; we use features from the second-last layer of CLIP instead of the last layer. Any image indicator tokens, e.g., special tokens enclosing visual tokens, are not used. We train the entire LLM instead of parameter-efficient fine-tuning. For in-depth ablations, we use a short training schedule (50k pre-training, 4k instruction tuning) with Vicuna-7B, CLIP ViT-L/14, and C-Abstractor with =144 visual tokens unless stated otherwise. For the final models, we adopt a long training schedule (200k pre-training, 10k instruction tuning). More details are in Appendix C.
5.2 Analysis on Locality-Enhanced Projector
| Projector | s/step | MME | MMB | SEED | Avg | |||||
| POS | SR | OL | PR | SR | IL | |||||
| B1 | Linear | 144 | - | Unavailable due to inflexiblity | - | |||||
| B2 | Resampler | 144 | 2.28 | 75.0 | 22.2 | 43.2 | 62.5 | 47.5 | 50.6 | 43.9 |
| B3 | C-Abstractor | 144 | 2.23 | 135.0 | 24.4 | 54.3 | 66.7 | 49.0 | 58.8 | 53.5 |
| B4 | Linear | 256 | 3.04 | 140.0 | 24.4 | 40.7 | 70.8 | 48.9 | 60.9 | 52.6 |
| B5 | Resampler | 256 | 3.12 | 73.3 | 24.4 | 37.0 | 79.2 | 44.4 | 51.8 | 45.6 |
| B6 | C-Abstractor | 256 | 3.07 | 136.7 | 26.7 | 55.6 | 75.0 | 52.7 | 59.3 | 56.3 |
To showcase the value of the proposed projector, we assess and compare both performance and efficiency against existing projectors in Table 3 using six spatial understanding tasks from MME, MMBench, and SEED-Bench. First, Resampler (B2, B5) shows poor performance due to its lack of consideration for local context preservation, despite being flexible to the number of visual tokens . Second, Linear projector is limited to =256 (B4) due to its inflexibility (B1), leading to intractable computational costs in high-resolution of larger . Third, for the same computational budget (=256), our C-Abstractor offer significantly improved performance compared to linear one (52.6 (B4) vs. 56.3 (B6)). Lastly, with fewer visual tokens (=144), our C-Abstractor demonstrate improved performance (+0.9 point) and greater efficiency (3.04 (B4) vs. 2.23 (B3) s/step). This improvement suggests our locality-enhanced projector excels at abstracting visual features where it integrates local contexts from neighboring features and provides context-enriched visual tokens, thus enabling our projectors to outperform linear counterparts even with fewer visual tokens.
@@ 5.3 Hidden Recipe for Visual Instruction Tuning
Table 4 shows a comprehensive ablation study to identify the individual impact of datasets on various multimodal benchmarks. First, we investigate the impact of dataset diversity within each task by leveraging only a single dataset for each task group (D1 vs. D2). The overall performance drop highlights the importance of the dataset diversity within each task. Second, we explore the impact of each task by sequentially excluding specific tasks (D1 vs. D3-8). This reveals that task diversity is crucial for learning how to handle a variety of tasks; each task improves the performance of relevant benchmarks, VQA (Open) MME, VQA (MC) MMB and SEED, and captioning and instruction-following data LLaVA. Third, we inspect the impact of using existing vision-language data (D9 vs. D10). Excluding such data leads to significant decreases in MME, MMB and SEED benchmarks. This suggests that rich knowledge in existing vision-language datasets enhances MLLM’s perception understanding or visual reasoning capabilities. In summary, these experiments emphasize the importance of diversity in both tasks and datasets within each task.
@@
| Method | LLM | Projector | Vision Encoder | Res. | MMB | MME | MME | SEED | LLaVA |
@@
Apart from datasets and instructization strategies, training recipes also incorporate several subtle yet crucial design choices, including the selection of features in vision encoder, LLMs, LLM training techniques, image indicators, pre-training and instruction tuning iterations. These recipes are detailed in Appendix D.
Final recipe.
In summary, our final recipe is summarized as 1) adopting flexible, locality-preserving C-Abstractor or D-Abstractor; 2) leveraging diverse datasets for various tasks (Table 4); 3) applying selected ablation options in Table 5 and Appendix D—the application of per-dataset balancing with hand-crafted tuning, fine-grained templates, and multi-turn interactions with deduplication.
@@ 5.4 Putting It Altogether
Comparison with existing MLLMs.
In Table 6, we compare our Honeybee, trained using the final recipe and a long training schedule, with other state-of-the-art MLLMs. Honeybee outperforms comparable 7B-scale MLLMs in all benchmarks, except for SEED. It is worth noting that competing methods like Qwen-VL [2] and LLaVA-1.5 [33] use larger vision encoders (e.g., ViT-bigG for Qwen-VL) or larger images (448 and 336) with more visual tokens (=256 and 576). In contrast, Honeybee employs ViT-L/14 with 224 resolution and 144 visual tokens striking a balance between performance and efficiency (Figure 8). For tasks requiring detailed visual understanding, such as SEED (see Appendix F), using larger images or more visual tokens can be beneficial. When the number of visual tokens is increased from 144 to 256, Honeybee achieves the best score in SEED (65.5) among 7B-scale LLMs, as shown in Table 7. When scaled up to 13B, Honeybee surpasses all previous methods in every benchmark. The detailed scores are available in Sec. G.1.
@@
In our final 7B and 13B models, we use 144 and 256 visual tokens (), respectively, balancing efficiency and performance. As indicated in Fig. 1 and Appendix A, increasing consistently improves performance. Our experiments, aligning in Honeybee with that of linear projector (Table 7), show performance enhancement at the cost of efficiency. Additional comparisons with previous methods are in Sec. G.2.
@@
Additional results.
@@ 6 Conclusion
The advent of visual instruction tuning has brought remarkable advances in MLLMs. Despite these strides, areas such as projector design and the approach in handling multifaceted data with instructization processes remain underexplored or unclear. Inspired by this, we identify the desirable but overlooked projector property, i.e., locality preservation, and propose the locality-enhanced projector that offers a preferable performance-efficiency balance. In addition, we provide extensive experiments to identify the impact of individual design choices in handling multifaceted instruction data, unveiling hidden recipes for high-performing MLLM development. Finally, Honeybee remarkably outperforms previous state-of-the-art methods on various benchmarks.
@@ B.1 Linear Projector
In the recent study, LLaVA (v1.5) [33] utilizes a 2-layer MLP instead of a single linear projection for enhancing the vision-language connector’s representation power. This approach led to an investigation of how varying the number of MLP layers impacts overall performance. As shown in Table 9, the 2-layer MLP-based projector marginally improves the overall performance compared to the linear projector. However, we observe a slight performance drop when further increasing the number of MLP layers (i.e., 6-layer MLP). We note that our C-Abstractor and D-Abstractor achieve better or comparable benchmark scores while using fewer visual tokens, indicating our projectors’ superiority regarding the balance of efficiency and performance.
@@ B.2 Resampler
As described in the main text, our design focuses on two principles: 1) flexibility in visual token counts, which is the key factor to the efficiency of MLLM, and 2) preservation of local context, which is critical for spatial understanding. Our first try is augmenting visual features with positional embeddings in the resampler framework, but it does not yield notable improvements (See Resampler in Table 9). This leads us to design two novel projectors, C-Abstractor and D-Abstractor.
@@ Appendix C Implementation Details
The detailed hyperparameters (HPs) are summarized in Table 11. Additionally, we utilize total six blocks in both C-Abstractor and D-Abstractor (i.e., for C-Abstractor and for D-Abstractor in Fig. 4). We use a single node with A100 80GB 8, employing deepspeed zero-2 [47] and flash-attention v2 [12] for all experiments, except for the pre-training of long schedule where we use multi-node setups.
l|l|m12cm
@@
A-OKVQA Human: Answer with the option’s letter from the given choices directly. {question} There are several options: {option} AI: {answer}